# Achieving 70 Gb/s Over A VCSEL-Based Optical Wireless Link Using A Multi-Mode Fiber-Coupled Receiver

### v1

@@ IV Conclusions
In this paper, we have demonstrated an experimental proof-of-concept for ultra-high data rate laser-based optical wireless links for future bandwidth-intensive applications. To this end, we employ a multi-aperture SM VCSEL operating at nm with an ultra-wide modulation bandwidth of GHz. We also employ a GHz MM fiber-optic photoreceiver to allow the VCSEL’s modulation bandwidth to be utilized by applying DCO-OFDM with adaptive subcarrier bit and power loading. The experimental results corroborate that the proposed system setup achieves a record-high data rate of Gb/s at a pre-FEC BER threshold of which is below the LTE target (i.e., ), while it attains Gb/s at a pre-FEC BER threshold of , over a link distance of cm. The link range and the received SNR can be increased using meticulous alignment procedures to maximize optical coupling efficiency at the MM fiber port of the receiver collimator. Furthermore, the GHz cutoff frequency of the oscilloscope is currently limiting the end-to-end performance of the system. Otherwise, considering the VCSEL bandwidth and the SNR response, this system can achieve beyod Gb/s. Therefore, more work is yet to be done to realize the full potential of the SM VCSEL for ultra-high speed OWC system design.

### v2

@@ IV Conclusions
In this paper, we have demonstrated an experimental proof-of-concept for ultra-high data rate laser-based optical wireless links for future bandwidth-intensive applications. To this end, we employ a SM MA VCSEL operating at nm with an ultra-wide modulation bandwidth of GHz. We also employ a GHz MM fiber-optic photoreceiver to allow the VCSEL’s modulation bandwidth to be utilized by applying DCO-OFDM with adaptive subcarrier bit and power loading. The experimental results corroborate that the proposed system setup achieves a record-high data rate of Gb/s at a pre-FEC BER of which is below the LTE target (i.e., ), while it attains Gb/s at a pre-FEC BER of , over a link distance of cm. The link range and the received SNR can be increased using meticulous alignment procedures to maximize optical coupling efficiency at the MMF port of the receiver collimator. Furthermore, the GHz cutoff frequency of the oscilloscope is currently limiting the end-to-end performance of the system. Otherwise, considering the VCSEL bandwidth and the SNR response, this system can reach data rates beyond Gb/s. Through extrapolating the SNR response in the absence of the hard cutoff frequency at GHz, it was verified that this system can potentially achieve Gb/s at a pre-FEC BER of . Therefore, more work is yet to be done to realize the full potential of the SM MA VCSEL for ultra-high speed OWC system design.
Appendix A Performance Bounds of MMF Under Dispersion
The maximum bandwidth and reachable distance of MMF links under dispersion are derived, assuming Gaussian impulse responses for both CD and MD [17]. The RMS delay spread due to CD is given by [17]:
where is the CD coefficient in ps/nmkm, is the RMS spectral width of the laser source in nm, and is the fiber length in km. The dB cutoff frequency for the CD-induced pulse broadening is obtained as [17]:
The dB cutoff frequency corresponding to the MD-induced pulse broadening is [27]:
where the EMB of MMF cables is typically given in MHzkm. The RMS delay spread associated with MD is then given by:
The total RMS delay spread under the joint effect of CD and MD can be expressed as [17]:
which leads to the combined dB cutoff frequency of:
By substituting (9) and (10) into (13), the maximum bandwidth supported by MMF for a running distance of is derived as:
The maximum reachable distance is determined by the upper bound of such that based on (14), with representing the signal bandwidth, resulting in: